# Unifying AI Tutor Evaluation: An Evaluation Taxonomy for Pedagogical Ability Assessment of LLM-Powered AI Tutors

### v1

@@ 7 Conclusion
This study provides a useful starting point for streamlining the evaluation of AI tutors. However, we acknowledge that there are certain limitations of this work, and addressing these limitations in the future is an important task.
@@
The proposed taxonomy primarily focuses on the task of the student mistake remediation in the domain of mathematics. We acknowledge that the proposed taxonomy will need to be verified on and likely adapted if applied to other tasks such as concept learning, and subjects other than mathematics. Nonetheless, we believe the proposed taxonomy, grounded in the learning sciences principles, will provide useful guidelines for future research.
@@
The current taxonomy and annotation scheme focus on the appropriateness of tutor responses. However, one of the limitations is that it does not consider the tutoring dialogues’ impact on the overall student learning. Specifically, the annotation pertains to the individual tutor turns within educational dialogues, which restricts our understanding of broader implications on student learning processes and learning gains, typically observed after a conversation concludes. We believe that the atomic tutor response evaluation at the utterance level, as presented in this study, should in the future be scaled up to the conversation level to better assess AI tutors’ pedagogical abilities.
@@
In this study, we limit the LLM-based evaluation to the Prometheus2 LLM as the critic, employing the evaluation prompt depicted in Figure 6. The results obtained with Prometheus2 are not encouraging, as detailed in Section 6. Future research should explore state-of-the-art and more powerful LLMs as critics and experiment with diverse prompt templates. We believe this preliminary study provides a foundational footprint for further investigations.
@@ Appendix A Evaluation Taxonomy, Annotation Labels, and Desiderata
Mistake identification Has the tutor identified a mistake in a student’s response? (1) Yes (2) To some extent (3) No Yes
@@
Coherence Is the tutor’s response logically consistent with the student’s previous response? (1) Yes (2) To some extent (3) No Yes
@@
Table 4: An overview of the proposed evaluation taxonomy, including associated annotation labels and desired expected labels. *For the guidance dimension, we provide further details for the labels: ‘Yes’ indicates that the guidance is correct and relevant to the mistake, ‘To some extent’ indicates that guidance is provided but is either partially/fully incorrect or incomplete, and ‘No’ indicates that no guidance has been provided.
@@
Table 5: Annotation correlation (AC) score between human and LLM-based judgements across different LLMs and evaluation dimensions on MRBench. The correlation scores are computed using Pearson’s correlation Sedgwick (2012). *For Novice, we have considered only 60 dialogues, whereas for Expert and all LLMs, all 192 instances were used.
Appendix B Prompt Template for Generating Responses from LLMs
Figure 2: The prompt template used to generate responses from the seven considered LLMs for both Bridge and MathDial datasets. The template is adapted from Wang et al. (2024a). The notable difference are: (1) the tutor in the Bridge dataset is for elementary school, whereas in MathDial it is for middle school and (2) The conversation topic is not available in MathDial.
@@ Appendix D Benchmark Statistics
#Total annotations* 540x8 1056x8 1596x8
@@ Appendix E Prompt Template for LLM-based Evaluation
Figure 6 illustrates the prompt template we have adapted for the evaluation with the Prometheus2 LLM Kim et al. (2024). The template is based on the insights drawn from the model’s official guidelines. [^3] https://github.com/prometheus-eval/prometheus-eval
Figure 6: Prompt template for evaluation with the Prometheus2 LLM Kim et al. (2024).

### v2

@@ 7 Conclusion
We believe that this study provides a useful starting point for streamlining the evaluation of AI tutors. However, we acknowledge that there are certain limitations of this work, and addressing these limitations in the future is an important task.
Establishing the relationships between evaluation dimensions
This study evaluates tutor response quality across the proposed eight dimensions independently. However, in practice, these dimensions may be inherently interrelated and may influence one another. A comprehensive investigation of these interdependencies can facilitate more effective modeling and ranking of tutor responses according to their quality at the dialogue level. The annotations provided in MRBench serve as a foundational resource for future research in this direction.
@@
The proposed taxonomy primarily focuses on the task of the student mistake remediation in the domain of mathematics. We acknowledge that the proposed taxonomy will need to be verified on and likely adapted if applied to other tasks such as concept learning, and to subjects other than mathematics. However, we believe that the proposed taxonomy, grounded in the learning sciences principles, will provide useful guidelines for future research.
@@
The current taxonomy and annotation scheme focus on the appropriateness of the tutor responses. However, one of the limitations is that it does not consider the tutoring dialogues’ impact on the overall student learning. Specifically, the annotation pertains to the individual tutor turns within educational dialogues, which restricts our understanding of broader implications on student learning processes and learning gains, typically observed after a conversation concludes. We believe that the atomic tutor response evaluation at the utterance level, as presented in this study, should in the future be scaled up to the conversation level to better assess AI tutors’ pedagogical abilities.
@@
In this study, we limit the LLM-based evaluation to two LLMs as critics, using the evaluation prompt presented in Figure 6. The results obtained with these LLMs are not encouraging, as detailed in Section 6. Future research should explore state-of-the-art and more powerful LLMs as critics and experiment with diverse prompt templates. At the same time, we believe that this preliminary study provides a basis and a benchmark for further investigation.
@@ Appendix A Evaluation Taxonomy, Annotation Labels, and Desiderata
Completeness of the evaluation taxonomy:
Through an iterative analysis of the taxonomy, we identify eight dimensions that comprehensively assess tutor response quality in the context of mistake remediation. However, other educational settings, particularly those involving tutorial dialogues beyond mistake remediation, may require modifications, as discussed in the limitations section. To establish a robust framework, we initially considered additional dimensions such as grammaticality and empathy, among others. However, our validation pilot study (see Section 4.2) confirmed that the selected eight dimensions are both necessary and sufficient for evaluating tutor response quality in dialogues aimed at mistake remediation.
@@
Mistake identification Has the tutor identified/recognized a mistake in a student’s response? (1) Yes (2) To some extent (3) No Yes
@@
Coherence Is the tutor’s response logically consistent with the student’s previous responses? (1) Yes (2) To some extent (3) No Yes
@@
Table 4: An overview of the proposed evaluation taxonomy, including associated annotation labels and desired expected labels. *For the guidance dimension, we provide further details for the labels: ‘Yes’ indicates that guidance is correct and relevant to the mistake; ‘To some extent’ indicates that guidance is provided but is either partially/fully incorrect or incomplete; and ‘No’ indicates that no guidance has been provided.
@@
Table 5: Annotation correlation (AC) scores between human annotations and judgments from Prometheus2 as LLM critic across different tutors and evaluation dimensions on MRBench. The correlation scores are calculated using Pearson’s correlation Sedgwick (2012). *Only 60 dialogues were considered for Novice, whereas all 192 dialogues were considered for Expert and other tutors.
*Novice -0.42 0.06 -0.71 -0.80 0.17 -0.17 -0.77 0.14
Expert -0.03 -0.29 -0.17 -0.23 -0.10 -0.16 -0.49 -0.01
Phi3 -0.71 -0.67 -0.77 -0.73 -0.61 -0.41 -0.62 0.04
Llama-3.1-8B -0.08 -0.46 -0.17 0.09 -0.09 -0.23 -0.38 0.09
Gemini 0.06 0.12 -0.11 -0.27 -0.22 -0.09 -0.34 0.09
Sonnet -0.07 -0.17 -0.26 -0.21 -0.29 -0.08 -0.32 0.07
Mistral -0.16 -0.16 -0.16 -0.34 -0.27 -0.28 -0.17 0.07
GPT-4 -0.03 0.01 -0.13 -0.23 0.01 -0.05 -0.06 0.10
Llama-3.1-405B -0.01 -0.02 -0.01 -0.07 0.00 0.02 -0.06 0.09
Table 6: Annotation correlation (AC) scores between human annotations and judgments from Llama-3.1-8B as LLM critic across different tutors and evaluation dimensions on MRBench. The correlation scores are calculated using Pearson’s correlation Sedgwick (2012). *Only 60 dialogues were considered for Novice, whereas all 192 dialogues were considered for Expert and other tutors.
Appendix B Prompt Template for Generating LLM Responses
Figure 2: The prompt template used to generate responses from the seven considered LLMs for both Bridge and MathDial datasets. The template is adapted from Wang et al. (2024a). The notable differences are: (1) The problems covered in the Bridge dataset are at the elementary school level, whereas those in MathDial are at the middle school level; and (2) The conversation topic is not provided in MathDial.
@@ Appendix D Benchmark Statistics
#Total annotations* 540$\times$8 1056$\times$8 1596$\times$8
@@ Appendix E Prompt Template for LLM-based Evaluation
Figure 6 illustrates the prompt template we have adapted for the evaluation with LLMs as critics Kim et al. (2024). The template is based on the insights drawn from the Prometheus2 model’s official guidelines. [^3] https://github.com/prometheus-eval/prometheus-eval
Figure 6: Prompt template for evaluation with LLMs as critics Kim et al. (2024).